OpenStudy (buttermequeasy):

Choose the correct procedure used to graph the equation 2x + 4y = 24 using the intercepts method. Solve the equation for y to get a y-intercept of (0, 6) and a slope of -1/2. Plot the y-intercept, count the slope to get another point and draw the line that passes through the two points. Let x equal zero and solve for y to find a y-intercept of (0, 6). Let y equal zero and solve for x to find an x-intercept of (12, 0). Plot the two points and draw a line passing through the two points. Choose two different values; plug them into the equation to find the corresponding y-values. Plot the

OpenStudy (buttermequeasy):

help please!!!

OpenStudy (anonymous):

Todos os três procedimentos são corretos. \[2x+4y=24 \rightarrow x+2y=12 \rightarrow 2y=12-x \rightarrow y=(12-x)/2\] That is a linear function with slope equal to -1 / 2 (the number that comes in front of x) To check each option, simply replace the values ​​as he suggests. You find the same values ​​that are indicated in the exercise.